1. The airbag is a safety device used in cars to cushion the passenger during a crash. It involves the following chemical reaction which is triggered by an impact:

2 NaN3 (s) → 2 Na(S) + 3N2 (g)

       Calculate the volume of N2 gas (measured in STP) that can be produced from 130.0 g of NaN3.

 

  1. How many liters of O2 (g) at STP is needed yo completely burn 20.0 L of butane (C4H10), with this reaction?

2 C4H10 (g) + 13 O2 (g) → 8 CO2 (g)  + 10 H2(l)

 

  1. Acetylene is formed by the reaction of water with calcium carbide, according to the following equation:

CaC2 (s) + 2H2(l) + Ca(OH)2 (aq) → C2H2 (g)

   How many grams of CaC2 would be needed to produce 10.0 L (at STP) of acetylene


1
Expert's answer
2022-01-12T13:03:03-0500

Number of moles NaN3= 130/65= 2

Mole ratio NaN3:N2= 2:3

= 2×3/2= 3moles

At STP, 1 mole = 22.4L

3 × 22.4L = 67.2L